OpenStudy (anonymous):

all of the following are ways to write 25 percent of N except a. 0.25N b. 25N/100 c. 1/4N d. 25N

OpenStudy (callisto):

25 percent of N = ( 25/100 ) N => never greater than N So, what is the answer?

OpenStudy (anonymous):

0.25N

OpenStudy (anonymous):

0.25 = 25/100 = 1/4

OpenStudy (anonymous):

kk thanks nightwill

OpenStudy (anonymous):

Welcome ^^

OpenStudy (callisto):

There is only one answer what will give something greater than N ....